Custom Query (419 matches)

Filters
 
Or
 
  
 
Columns

Show under each result:


Results (58 - 60 of 419)

Ticket Resolution Summary Owner Reporter
#80 fixed GB2005/03, flight day 254a/2007, Harwood mggr mggr
Description

Data location: ~arsf/arsf_data/in_progress/2007/flight_data/uk/GB2005_03-2007_254a_Harwood

Data arrived from ARSF via SATA disk transfer on 29th Oct.

Scientific details: (details of 2005 applications not available at PML)

PI: Tim Malthus

Sensors:

  • ATM (delivered 30/Apr/2008)
  • CASI (delivered 20/Dec/2007)
  • Eagle (delivered 12/Mar/2008)
  • Hawk (delivered 30/Apr/2008)

Quicklook

#81 fixed Eagle radiometric calibration mismatch mggr mggr
Description

This ticket is spawned from #74 (support ticket for Rachel Gaulton, who found this problem). See that ticket for the initial work, summarised here.

Eagle's calibrated spectral profile on flights appears to differ substantially from CASI in terms of the measured radiance. The shape is the same but the numbers are different by a factor of approximately 4-8, depending on the pixel chosen.

Short history to date:

  • 16/Oct Initial problem report from Rachel (2006 dataset)
  • 18/Oct AKW recommended investigating on a known dataset
  • 22/Oct Picked 166/2007 boresight flight and confirmed we're also seeing this issue
  • 29/Oct Demonstrated to Gary and agreed routes for investigation

Original L0 and L1b tests

#82 fixed GB2004/03, flight day 171a/2007, Huntingdon/Cambridge mggr mggr
Description

Data location: ~arsf/arsf_data/in_progress/2007/flight_data/uk/GB2004_03-2007_171a_Huntingdon_Cambs Data arrived from ARSF via second SATA disk transfer at the start of Sept.

Scientific details:

  • (2004 applications not held at PML)

PI: Shelley Hinsley

This project was originally bundled together with #58 (two jobs done in one flight, but were kept as a single logsheet, hence the confusion). Split out into two jobs on 13/Nov. Should be fully processed.

Note: See TracQuery for help on using queries.